OpenStudy (anonymous):

Choose the slope-intercept equation of the line that passes through the point (-2, 2) and is parallel to y = 4x + 7.

OpenStudy (anonymous):

the slope of the line is 4 since it is parallel to the line y=4x+7, hence they have the same slope and since it passes through (-2,2) we'll have, y-2=4(y+2) y-2=4y+8 y=4y+10 <---answer